# Computation of the semiclassical outflux emerging from a collapsing spherical null shell

### v1

@@ Acknowledgements.
We are grateful to Paul Anderson, Jochen Zahn, Stefan Hollands, Marc Casals, Adrian Ottewill and Maria Alberti for interesting discussions and helpful feedback. We are especially thankful to Adam Levi for providing the digital version of the Unruh-state RSET data of Ref. (LeviRSET:2017, ) used for the preparation of Figure 2 in the Appendix. NZ acknowledges the generous support of Fulbright Israel.
.1 Appendix
As mentioned in the Discussion, the function $F^{+}\left(r_{0}\right)$ – representing the outflux on the Schwarzschild side of the shell as a function of $r_{0}$ (namely the area coordinate $r$ on the shell) – attains a maximum value $\approx 1.031\times 10^{-4}\hbar M^{-2}$ at $r_{0}^{\text{max}}=(10/3)M$ that exceeds the known (Elster:1983, ) Hawking radiation outflux $F_{H}\approx 7.44\times 10^{-5}\hbar M^{-2}$. This fact may suggest a nonmonotonic dependence of the outflux $F$ on $r$ along the outgoing null geodesic emerging from the shell at $r_{0}^{\text{max}}$  [^5] Note, however, that there is another possibility: The outflux at FNI, namely $\lim_{r\to\infty}4\pi r^{2}\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$, may in principle be a nonmonotonic function of $u$ (such that this FNI-outflux decreases with $u$ at some range of $u$ values later than the outgoing null geodesic that emanates from the shell at $r_{0}^{\text{max}}$). (as well as along its neighboring outgoing null geodesics). To put this suggested behavior in a wider context, it may be useful to compare it with the $r$-dependence of the outflux in the Unruh state in Schwarzschild, denoted $F^{U}\left(r\right)\equiv 4\pi r^{2}\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}^{U}\left(r\right)$ (here, the superscript "$U"$ denotes the Unruh state). Due to the stationarity in this case, $F^{U}$ is a function of $r$ alone, satisfying $\lim_{r\to\infty}F^{U}(r)=F_{H}$.
Using the results for $\left\langle\hat{T}_{t}^{t}\right\rangle_{\text{ren}}^{U}$, $\left\langle\hat{T}_{r}^{r}\right\rangle_{\text{ren}}^{U}$ and $\left\langle\hat{T}_{t}^{r}\right\rangle_{\text{ren}}^{U}$ of Ref. (LeviRSET:2017, ) (see, in particular, Fig. 4 therein), we construct $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}^{U}\left(r\right)$ and then $F^{U}\left(r\right)$, and explore the latter’s dependence on $r$. As expected from regularity, we find that it behaves as $\left(1-2M/r\right)^{2}$ near the EH (similar to $F^{+}\left(r\right)$, though with a different prefactor). Then, as $r$ increases outwards, the Unruh outflux does not grow monotonically toward $F_{H}$. Instead, it initially increases to a maximum at $r\approx 4.55M$, see Fig. 2. Then, it gradually decreases toward $F_{H}$ as $r\to\infty$.
Given this behavior in the Unruh state, it is not too surprising if the outflux along outgoing null geodesics in the collapsing-shell spacetime exhibits a similar nonmonotonic trend.
Figure 2: The black dotted curve represents the Unruh outflux, $F^{U}=4\pi r^{2}\left\langle T_{uu}\right\rangle_{\text{ren}}^{U}$, as a function of $r$, asymptotically approaching $F_{H}$ which is indicated by the horizontal dashed line. The horizontal blue line marks the maximum value of $F^{+}\left(r_{0}\right)$, the outflux emerging from the shell. Notably, this value is greater than $F_{H}$ but still smaller than the peak of $F^{U}\left(r\right)$.

### v3

@@ Acknowledgements.
We are grateful to Paul Anderson, Jochen Zahn, Stefan Hollands, Marc Casals, Adrian Ottewill and Maria Alberti for interesting discussions and helpful feedback. We are especially thankful to Adam Levi for providing the digital version of the Unruh-state RSET data of Ref. (LeviRSET:2017, ) used for the preparation of Fig. [1] in the Supplemental Material. N.Z. acknowledges the generous support of Fulbright Israel.
?refname?
  • (1) S. W. Hawking, Black hole explosions?, Nature 248, 30 (1974).
  • (2) S. W. Hawking, Particle creation by black holes, Comm. Math. Phys. 43, 199 (1975).
  • (3) W. G. Unruh, Origin of the particles in black-hole evaporation, Phys. Rev. D 15, 365 (1977).
  • (4) P. Chen, W. G. Unruh, C. H. Wu, and D. H. Yeom, Pre-hawking radiation cannot prevent the formation of apparent horizon, Phys. Rev. D 97, 064045 (2018).
  • (5) P. C. W. Davies, On the Origin of Black Hole Evaporation Radiation, Proc. R. Soc. A 351, 129 (1976).
  • (6) U. H. Gerlach, The mechanism of blackbody radiation from an incipient black hole, Phys. Rev. D 14, 1479 (1976).
  • (7) D. G. Boulware, Hawking radiation and thin shells, Phys. Rev. D 13, 2169 (1976).
  • (8) H. Kawai, Y. Matsuo, and Y. Yokokura, A Self-consistent Model of the Black Hole Evaporation, Int. J. Mod. Phys. A 28, 1350050 (2013).
  • (9) P. M. Ho, Comment on Self-Consistent Model of Black Hole Formation and Evaporation, J. High Energy Phys. 08 (2015) 096.
  • (10) H. Kawai and Y. Yokokura, Interior of Black Holes and Information Recovery, Phys. Rev. D 93, 044011 (2016).
  • (11) P. M. Ho, The Absence of Horizon in Black-Hole Formation, Nucl. Phys. B909, 394 (2016).
  • (12) D. Maskalaniec and B. Sikorski, Hawking radiation far away from the event horizon, arXiv:2409.11021v1 [gr-qc].
  • (13) L. Mersini-Houghton, Backreaction of Hawking Radiation on a Gravitationally Collapsing Star I: Black Holes?, Phys. Lett. B 738, 61 (2014).
  • (14) L. Mersini-Houghton, Back-reaction of the Hawking radiation flux on a gravitationally collapsing star II, arXiv:1409.1837 [hep-th].
  • (15) V. Baccetti, R. B. Mann, and D. R. Terno, Role of evaporation in gravitational collapse, Classical Quantum Gravity 35, 185005 (2018).
  • (16) V. Baccetti, R. B. Mann, and D. R. Terno, Horizon avoidance in spherically-symmetric collapse, arXiv:1703.09369 [gr-qc].
  • (17) V. Baccetti, R. B. Mann, and D. R. Terno, Do event horizons exist?, Int. J. Mod. Phys. D 26, 743008 (2017).
  • (18) P. R. Anderson, S. G. Siahmazgi, R. D. Clark, and A. Fabbri, Method to compute the stress-energy tensor for a quantized scalar field when a black hole forms from the collapse of a null shell, Phys. Rev. D 102, 125035 (2020).
  • (19) S. G. Siahmazgi, P. R. Anderson, R. D. Clark, and A. Fabbri, Stress-energy Tensor for a Quantized Scalar Field in a Four-Dimensional Black Hole Spacetime that Forms From the Collapse of a Null Shell, in The Sixteenth Marcel Grossmann Meeting on Recent Developments in Theoretical and Experimental General Relativity, Astrophysics and Relativistic Field Theories: Proceedings of the MG16 Meeting on General Relativity (2021), pp. 1265-1274.
  • (20) S. G. Siahmazgi, P. R. Anderson, and A. Fabbri, Late-time behaviors of scalar field modes for a collapsing null shell spacetime and for the Unruh state in Schwarzschild spacetime, arXiv:2502.12910 [gr-qc].
  • (21) S. Massar and R. Parentani, From vacuum fluctuations to radiation. II. Black holes, Phys. Rev. D 54, 7444 (1996).
  • (22) A. Fabbri and J. Navarro-Salas, Modeling Black Hole Evaporation (Imperial College Press, London, 2005).
  • (23) B. S. DeWitt, Dynamical Theory of Groups and Fields (Gordon and Breach, New York, 1965).
  • (24) A. Levi and A. Ori, Mode-sum regularization of $\langle{\phi}^{2}\rangle$ in the angular-splitting method, Phys. Rev. D 94, 044054 (2016).
  • (25) P. Candelas, Vacuum polarization in Schwarzschild spacetime, Phys. Rev. D 21, 2185 (1980).
  • (26) A. Levi, Renormalized stress-energy tensor for stationary black holes, Phys. Rev. D 95, 025007 (2017).
  • (27) N. Zilberman, A. Levi, and A. Ori, Quantum Fluxes at the Inner Horizon of a Spherical Charged Black Hole, Phys. Rev. Lett. 124, 171302 (2020).
  • (28) T. Elster, Vacuum polarization near a black hole creating particles, Phys. Lett. A 94, 205 (1983).
  • (29) A. Levi and A. Ori, Pragmatic mode-sum regularization method for semiclassical black-hole spacetimes, Phys. Rev. D 91, 104028 (2015).
  • (30) A. Levi and A. Ori, Versatile Method for Renormalized Stress-Energy Computation in Black-Hole Spacetimes, Phys. Rev. Lett. 117, 231101 (2016).
  • (31) N. Zilberman, M. Casals, A. Ori, and A. C. Ottewill, Quantum fluxes at the inner horizon of a spinning black hole, Phys. Rev. Lett. 129, 261102 (2022).
  • (32) A. Levi, E. Eilon, A. Ori, and M. van de Meent, Renormalized Stress-Energy Tensor of an Evaporating Spinning Black Hole, Phys. Rev. Lett. 118, 141102 (2017).
  • (33) A. Lanir, A. Levi, and A. Ori, Mode-sum renormalization of $\langle\hat{\Phi}^{2}\rangle$ for a quantum scalar field inside a Schwarzschild black hole, Phys. Rev. D 98, 084017 (2018).
  • (34) N. Zilberman, M. Casals, A. Levi, A. Ori, and A. C. Ottewill, Computation of $\left\langle\Phi^{2}\right\rangle$ and quantum fluxes at the polar interior of a spinning black hole, Phys. Rev. D 111, 124054 (2025).
  • (35) W. A. Hiscock, Models of evaporating black holes. I, Phys. Rev. D 23, 2813 (1981).
  • (36) P. C. W. Davies, Quantum Vacuum Stress Without Regularization in Two-Dimensional Space-Time, Proc. R. Soc. A 354, 1679 (1977).
  • (37) A. Lanir, A. Levi, A. Ori, and O. Sela, Two-point function of a quantum scalar field in the interior region of a Reissner-Nordstrom black hole, Phys. Rev. D 97, 024033 (2018).
  • (38) See Supplemental Material for the $r$-dependence of the outflux in the Unruh state in Schwarzschild, based on data from Ref. (LeviRSET:2017, ). The Supplemental Material also includes Ref. (Elster:1983, ).